OpenStudy (boxman61):

Do not know the process to solve this answer 2.05*10^-18= 2.18*10^-18 (1/Nf^2 - 1/Ni^2)

OpenStudy (issimplcalcus):

What are you solving for?

OpenStudy (boxman61):

SO this is a question from Chemistry where we have a final reading and an initial reading. I am trying to find the final reading (Nf)

OpenStudy (boxman61):

\[2.05*10^-18 = -2.18*10^-18 (\frac{ 1 }{ n _{f}^2 }-\frac{ 1 }{ n _{i}^2 })\]

OpenStudy (issimplcalcus):

You can find Nf in terms of Ni just by algebra.

OpenStudy (boxman61):

\[-0.9404= (\frac{ 1 }{ n _{f}^2 }-\frac{ 1 }{ n _{i}^2 })\]

OpenStudy (boxman61):

This is where i'm stuck. My teacher has -0.9404 = (1/nf^2 - 1/1^2) then 0.0596 = 1/nf^2 But I dont know how she deducted the 1/1^2